# Requirements are All You Need: From Requirements to Code with LLMs

### v1

@@ I Introduction
In the realm of Software Engineering, Requirements Engineering (RE) is pivotal, serving as the foundation upon which all subsequent development stages are built. Requirements are traditionally documented in natural language [1], varying from structured formats like use cases to informal descriptions. The textual nature of requirements inspires researchers to apply natural language processing technologies for various requirements engineering tasks [2]. The advent of large language models (LLMs) such as ChatGPT offers a promising avenue for assisting requirements engineering processes. These language models, powered by deep learning algorithms and gigantic training corpus, exhibit an unprecedented ability to understand, analyze, and generate natural language.
@@
The second assumption is that LLMs need software engineers’ collaboration to be effective. Instead of letting a LLMs take full control, they should behave like an adviser for the software engineers. The LLMs and the software engineers ought to collaboratively find out the design details that will be used in the implementation of the software system. By automating parts of the analysis, design, and coding process, LLMs allow software engineers to focus on more complex and creative aspects of software development.
@@
  • We introduce a tailored LLM for generating source code based on the provided use case document. The LLM is fine-tuned on the knowledge, instructions, and heuristics pertinent to software development process, requirements analysis, object-oriented design, and test-driven development. The LLM is available at https://chat.openai.com/g/g-bahoiKzkB-software-engineer-gpt.
@@
The structure of this paper is organized in the following manner: Section II examines the existing literature on generating code with LLMs. Section III elaborates on the design specifics of our custom-developed LLM. Following that, Section IV offers a case study that exemplifies the model’s capability in comprehending requirements and producing code that is both syntactically and semantically accurate. Section V explores the implications of LLMs within the field of software engineering and outlines the directions for future research. Finally, Section VI concludes the paper.
@@ II Related Work
III A Tailored LLM for Software Engineering
In this section, we explore the design of our tailored LLM crafted to support software engineers in generating code from high-level requirements. To clearly depict the process of engagement with our customized LLM, we employ a UML activity diagram, which outlines the workflow for interaction between the software engineers and the LLM (refer to Fig. 1).
Figure 1: Human software engineer’s interaction with the tailored GPT model.
III-A Software Requirements Documents
Without a foundation of high-quality requirements as input, even the most performant LLM can be rendered useless. Instead of typing a few informal itemized requirements to an LLM, we require the software development team to upload to the LLM structured requirements in the format of project glossary, vision and scope, and use cases. The time and effort invested in writing high-quality software requirements pays off significantly, especially in the era of LLMs.
III-B A Progressive Prompting Approach
@@
This progressive prompting approach is in line with both the waterfall software development process and the requirements abstraction hierarchy (Fig. 2). By breaking down the complex software development task into smaller and more specific subtasks, this approach generates a “chain of thought” [5], making the LLM more efficient and effective to perform complex tasks in the software development. The intermediate artifacts like functional requirements and object-oriented design are valuable to the software development team since they give the human engineers an opportunity to understand and verify the LLM’s chain of thought and provide feedback. Those artifacts also help verify in case of hallucination. This will also have the benefit of requirement traceability [6].
@@
III-C Instructions and Knowledge Base for the Tailored LLM
@@
IV Case Study
To demonstrate LLMs’ capability of generating code from requirements, we chose a medium-sized web project called “SuperFrog Scheduler”. SuperFrog is the iconic horned frog mascot of Texas Christian University (TCU), known for its participation in numerous events annually, including sports, weddings, charity functions, and media appearances. To facilitate the scheduling of SuperFrog’s appearances, a request management system is necessary, involving submission by customers, review by the Spirit Director, and acceptance by student performers embodying SuperFrog. A development team, consisting of five software developers, commenced work on the SuperFrog Scheduler in October 2018 and successfully completed the project in May 2019. During the requirements gathering phase, the team identified and documented 26 user-goal level use cases, which are accessible via Google Docs. For the purposes of this preview paper, our focus is restricted to the implementation of Use Case 18. Use Case 18 describes the process where the Spirit Director generates TCU Honorarium Payment Request Forms for SuperFrog Students who have completed appearance requests. The system facilitates the creation of these forms by providing the necessary details for each eligible appearance. This process aims to ensure SuperFrog Student workers are compensated for their services in a timely and efficient manner.
@@ IV Case Study
IV-A Uploading Software Requirements
We initiated the process by uploading the project glossary, vision and scope, and use cases to the tailored GPT. These requirements documents provide critical context and background information to subsequent tasks needed. The GPT then summarized the contents of the three documents (see the conversation excerpt in Fig. 3). The GPT prompted us to pick one use case to start.
@@ IV-A Uploading Software Requirements
IV-B Deriving Functional Requirements
Use cases describe tasks that users will need to perform with the system, or user-system interactions that will result in a valuable outcome for some stakeholder. We instructed the GPT to derive necessary functional requirements that must be implemented to enable each use case. As a proof of concept, we only demonstrated the design of Use Case 18, which pertains to the generation of Honorarium (Payment for Services) request forms for student workers (see the derived functional requirements in the conversation excerpt in Fig. 4). Then, the GPT asked us to proceed with creating an object-oriented design for this use case.
@@ IV-B Deriving Functional Requirements
IV-C Generating an Object-Oriented Design
@@
IV-D Generating Test Cases
In this step (see Fig. 6), the tailored GPT was instructed to generate one unit test case for the generateForm method in the class FormGenerator. We also specified the technologies required. It turned out that again we needed to clarify some misunderstandings (detailed conversion is not shown).
@@ IV-D Generating Test Cases
IV-E Generating Implementation Code
In the end, based on the intermediate artifacts or “chain of thoughts” generated in the previous steps, we believe that the GPT has enough context and information to move to the final step. We then instructed the GPT to generate the implementation code for the generateForm in the class FormGenerator (see Fig. 7).
@@ V-A Requirements are All You Need
V-B Impact on Software Engineering Education
LLMs have transformed the landscape of programming, showcasing exceptional abilities in code generation and problem-solving. This technological advancement has undeniably increased efficiency and productivity within software development. However, the question arises: Is teaching traditional designing, programming, and testing skills to students still necessary? The unequivocal answer is yes. Despite the prowess of LLMs in handling a variety of programming tasks, their limitations cannot be overlooked. These models sometimes misunderstand us and produce code with errors or security vulnerabilities, underscoring the need for human oversight and validation. It is imperative for software engineering education to evolve, teaching students to critically assess AI-generated designs and code for accuracy, efficiency, maintainability, and security.
The emergence of LLMs, such as ChatGPT, and tools like GitHub Copilot, signals a pivotal shift in software engineering education. To stay ahead, educators must not only embrace the LLMs themselves but also guide students on how to effectively leverage LLMs. This entails a significant pedagogical shift towards enhancing soft skills. Students should be trained to excel in communication with clients, the art of drafting high-quality software requirements, and conducting thorough design and code reviews. Such skills are becoming increasingly vital as the role of manual coding, particularly for minor features, diminishes.
V-C LLMs Redefine Agile
@@
V-D Limitations and Future Work
The proof-of-concept case study showcased the significant potential of leveraging LLMs in software development, particularly in the context of use-case-driven development for web projects. However, further testing across a broader spectrum of software projects—including real-time systems, AI-driven applications, and games—is essential to fully understand the applicability and limitations of LLMs in these varied contexts. Additionally, exploring different requirements specification and design methodologies is crucial to assess the adaptability and effectiveness of LLMs in diverse software engineering paradigms. While a tailored version of ChatGPT has been developed and made publicly available, utilizing the default user interface of ChatGPT for software development presents practical challenges. Specifically, software engineers may find it cumbersome to navigate through previously generated designs or code snippets due to the need to scroll extensively. To address this issue, we envisage the creation of a user-friendly web application in future work. This application will aim to streamline the process, offering a more intuitive and visual approach to translating requirements into tests and code.

### v2

@@ I Introduction
In the realm of Software Engineering, Requirements Engineering (RE) is pivotal, serving as the foundation upon which all subsequent development stages are built. Requirements are traditionally documented in natural language [1], varying from structured formats like use cases to informal descriptions. The textual nature of requirements inspires researchers to apply natural language processing technologies for various requirements engineering tasks [2]. The advent of large language models (LLMs) such as ChatGPT offers a promising avenue for assisting requirements engineering processes. These language models, powered by deep learning algorithms and a gigantic training corpus, exhibit an unprecedented ability to understand, analyze, and generate natural language.
@@
The second assumption is that LLMs need software engineers’ collaboration to be effective. Instead of letting an LLM take full control, they should behave like a consultant for the software engineers. The LLMs and the software engineers ought to collaboratively find out the design details that will be used in the implementation of the software system. By automating parts of the analysis, design, and coding process, LLMs allow software engineers to focus on more complex and creative aspects of software development.
@@
  • We introduced a tailored LLM for generating source code based on the provided software requirements. The LLM is fine-tuned on the knowledge, instructions, and heuristics pertinent to the software development process, requirements analysis, object-oriented design, and test-driven development. The LLM is available at https://chat.openai.com/g/g-bahoiKzkB-software-engineer-gpt.
@@
The structure of this paper is organized in the following manner: Section II examines the existing literature on generating code with LLMs. Section III presents a detailed description of our research method, outlining the systematic approach we have developed to enhance the application of LLMs in software engineering. Section IV elaborates on the design specifics of our custom-developed LLM. Following that, Section V offers a case study that exemplifies the model’s capability in comprehending requirements and producing code that is both syntactically and semantically accurate. Section VI explores the implications of LLMs within the field of software engineering and outlines the directions for future research. Finally, Section VII concludes the paper.
@@ II Related Work
III Research Method
The review of existing literature reveals a significant gap in the application of LLMs for the direct generation of executable software code from natural language requirements. To bridge this gap, our research introduces a novel method that not only enhances the LLM’s understanding of software requirements but also iteratively refines its output to ensure that the generated code meets the semantic expectations. This methodological innovation is vital for advancing the use of LLMs in software development, particularly in environments where accuracy and alignment with business goals are paramount.
III-A Research Objectives
Our research aims to develop and evaluate a tailored LLM that can effectively translate detailed software requirements into high-quality, functional code. The primary objective of this research is to demonstrate the feasibility of using LLMs for end-to-end software development tasks, thereby minimizing human intervention in the coding process while ensuring that the outputs are practically viable and aligned with user intentions.
III-B Research Design
This research follows a phased approach:
  • Phase 1: Model Selection and Customization. We begin by selecting ChatGPT as the foundational LLM due to its advanced capabilities in understanding and generating natural language. The LLM is then customized to suit the specific needs of software development, integrating software engineering knowledge and instructions that enhance its ability to process and translate software requirements into functional code.
  • Phase 2: Data Preparation. Software requirements will be collected with the intent to represent the diversity of real-world software development projects as accurately as possible. These will include both industry-standard specifications and hypothetical scenarios designed to test the model’s capabilities across different scenarios.
  • Phase 3: Evaluation and Refinement. The tailored LLM’s effectiveness will be evaluated through a series of controlled tests, analyzing both the code quality and its alignment with the input requirements. The evaluation phase will involve comparing the LLM-generated code with expert-written code to assess accuracy, adherence to requirements, and practical usability. Both qualitative and quantitative measures will be employed to provide a thorough evaluation of the tailored LLM’s performance. Feedback from these evaluations will be used to further refine the model.
III-C Iterative Refinement Process
An iterative refinement process is integral to the methodology. After each evaluation phase, detailed feedback will be analyzed to identify any shortcomings of the tailored LLM. This feedback informs subsequent refinements to the model, enhancing the tailored LLM’s ability to generate more accurate and relevant code outputs.
IV A Tailored LLM for Software Engineering
In this section, we explore the design of our tailored LLM crafted to support software engineers in generating code from high-level requirements. To depict the process of engagement with our customized LLM, we employ a UML activity diagram, which outlines the workflow for interaction between the software engineers and the LLM (Fig. 1).
Figure 1: A human software engineer’s interaction with the tailored GPT model.
IV-A Software Requirements Documents
Without a foundation of high-quality requirements as input, even the most performant LLM can be rendered useless. Instead of typing a few informal itemized requirements to an LLM, we require the software development team to upload to the LLM structured requirements in the format of project glossary (Google Docs), vision and scope (Google Docs), and use cases (Google Docs).
  • Project Glossary: This document contains definitions of all key terms used within the project, ensuring clarity and consistency across all project communications and artifacts.
  • Vision and Scope: This outlines the overall objectives, the scope of the project, and the roles of different stakeholders, providing a comprehensive overview of the project’s goals and expected outcomes.
  • Use Cases: This document includes detailed descriptions of each use case, illustrating specific functionality that the software is expected to perform. These are particularly valuable for understanding the interactions between the system and its users, as well as between different components of the system.
The time and effort invested in writing high-quality software requirements pays off significantly, especially in the era of LLMs.
IV-B A Progressive Prompting Approach
@@
This progressive prompting approach is in line with both the waterfall software development process and the requirements abstraction hierarchy (Fig. 2). By breaking down the complex software development task into smaller and more specific subtasks, this approach generates a “chain of thought” [5], making the LLM more efficient and effective in performing complex tasks in software development. The intermediate artifacts like functional requirements and object-oriented design are valuable to the software development team since they allow the human engineers to understand and verify the LLM’s chain of thought and provide feedback. Those artifacts also help verify in case of hallucination. This will also have the benefit of requirement traceability [6].
@@
IV-C Instructions and Knowledge Base for the Tailored LLM
@@
V Case Study
To demonstrate LLMs’ capability of generating code from requirements, we chose a medium-sized web project called “SuperFrog Scheduler”. SuperFrog is the iconic horned frog mascot of Texas Christian University (TCU), known for its participation in numerous events annually, including sports, weddings, charity functions, and media appearances. To facilitate the scheduling of SuperFrog’s appearances, a request management system is necessary, involving submission by customers, review by the Spirit Director, and acceptance by student performers embodying SuperFrog. A development team, consisting of five software developers, commenced work on the SuperFrog Scheduler in October 2018 and completed the project in May 2019. During the requirements gathering phase, the team identified and documented 26 user-goal level use cases, which are available here (Google Docs). The “SuperFrog Scheduler” was selected for its complexity and relevance, involving multiple stakeholders and intricate scheduling requirements typical of real-world software projects. This case study not only provides detailed documentation for verifying the tailored LLM’s code generation capabilities but also showcases the model’s practical application in a dynamic educational environment. Its visibility and the educational context offer a unique opportunity to demonstrate the potential of LLMs in streamlining complex software development tasks while educating future software engineers on AI-driven development tools. For this preview paper, our focus is restricted to the implementation of Use Case 18 (Google Docs). Use Case 18 describes the process where the Spirit Director generates TCU Honorarium Payment Request Forms for SuperFrog Students who have completed appearance requests. The system facilitates the creation of these forms by providing the necessary details for each eligible appearance. This process aims to ensure SuperFrog Student workers are compensated for their services in a timely and efficient manner.
@@ V Case Study
To enhance the accessibility and reproducibility of our research, we have compiled all essential project artifacts into a replication package [^7] https://drive.google.com/drive/folders/1CFxMNblwOV-mthVcRRb2gTsFxNRO7nlz?usp=sharing. This package includes the project glossary, vision and scope, and detailed use cases, which are fundamental to understanding the development context and requirements of the “SuperFrog Scheduler”.
V-A Uploading Software Requirements
We initiated the process by uploading the project glossary, vision and scope, and use cases to the tailored GPT. These requirements documents provide critical context and background information for subsequent tasks needed. The GPT then summarized the contents of the three documents (see the conversation excerpt in Fig. 3). The GPT prompted us to pick one use case to start.
@@ V-A Uploading Software Requirements
V-B Deriving Functional Requirements
Use cases describe tasks that users will need to perform with the system, or user-system interactions that will result in a valuable outcome for some stakeholders. We instructed the GPT to derive the necessary functional requirements that must be implemented to enable each use case. As a proof of concept, we only demonstrated the design of Use Case 18, which pertains to the generation of Honorarium (Payment for Services) request forms for student workers (see the derived functional requirements in the conversation excerpt in Fig. 4). Then, the GPT asked us to proceed with creating an object-oriented design for this use case.
@@ V-B Deriving Functional Requirements
V-C Generating an Object-Oriented Design
@@
V-D Generating Test Cases
In this step (Fig. 6), the tailored GPT was instructed to generate one unit test case for the generateForm method in the class FormGenerator. We also specified the technologies required. It turned out that again we needed to clarify some misunderstandings (detailed conversion is not shown).
@@ V-D Generating Test Cases
V-E Generating Implementation Code
In the end, based on the intermediate artifacts or “chain of thoughts” generated in the previous steps, we believe that the GPT has enough context and information to move to the final step. We then instructed the GPT to generate the implementation code for the generateForm in the class FormGenerator (Fig. 7).
@@ VI-A Requirements are All You Need
VI-B Impact on Software Engineering Education
LLMs have transformed the landscape of programming, showcasing exceptional abilities in code generation and problem-solving. This technological advancement has undeniably increased efficiency and productivity within software development. However, the question arises: Is teaching traditional designing, programming, and testing skills to students still necessary? The unequivocal answer is yes. Despite the prowess of LLMs in handling a variety of programming tasks, their limitations cannot be overlooked. These models sometimes misunderstand us and produce code that is not aligned with our intent, underscoring the need for human oversight and validation. It is imperative for software engineering education to evolve, teaching students to critically assess AI-generated designs and code for accuracy, efficiency, maintainability, and security.
The emergence of LLMs, such as ChatGPT, and tools like GitHub Copilot and Devin AI, signals a pivotal shift in software engineering education. To stay ahead, educators must not only embrace the LLMs themselves but also guide students on how to effectively leverage LLMs. This entails a significant pedagogical shift towards enhancing soft skills. Students should be trained to excel in communication with various stakeholders, the art of drafting high-quality software requirements, and conducting thorough design and code reviews. Such skills are becoming increasingly vital as the role of manual coding, particularly for minor features, diminishes.
VI-C LLMs Redefine Agile
@@
VI-D Limitations and Future Work
The proof-of-concept case study showcased the significant potential of leveraging LLMs in software development, particularly in the context of use-case-driven development for web projects. However, further testing across a broader spectrum of software projects—including real-time systems, AI-driven applications, and games—is essential to fully understand the applicability and limitations of LLMs in these varied contexts. Additionally, exploring different requirements specifications and design methodologies is crucial to assess the adaptability and effectiveness of LLMs in diverse software engineering paradigms. While a tailored version of ChatGPT has been developed and made publicly available, utilizing the default user interface of ChatGPT for software development presents practical challenges. Specifically, software engineers may find it cumbersome to navigate through previously generated designs or code snippets due to the need to scroll extensively. To address this issue, we envisage the creation of a user-friendly web application in future work. This application will aim to streamline the process, offering a more intuitive and visual approach to translating requirements into tests and code.